# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ V Proof of Lemma 3
$\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}}{1-\gamma}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\upsilon\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+4\alpha^{2}\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}.$
@@
${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\text{ and }{\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}.$
@@
By $\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|=\frac{2\chi}{1-\lambda_{2}}$ and $L$-smoothness of $f_{i}$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 4L^{2}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{1-\lambda_{2}}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (52)
@@
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 16L^{2}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{1-\lambda_{2}}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (53)
@@
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $(16L^{2}+\frac{16\alpha^{2}L^{4}\chi}{1-\lambda_{2}})\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{2}L^{2}\chi}{1-\lambda_{2}}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{2}L^{2}\sigma^{2}\chi}{1-\lambda_{2}}.$ (54)
Thus, combining (V) and (54), it holds that
$\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=$ $\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$
$+\frac{8n\alpha^{4}L^{2}\chi}{(1-\gamma)(1-\lambda_{2})}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{4\alpha^{4}L^{2}\sigma^{2}\chi}{(1-\gamma)(1-\lambda_{2})}+4n\alpha^{2}\sigma^{2}.$ (55)
Then, we bound $\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$. Using $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, $\hat{{\bf{Z}}}^{t}={\bf{X}}^{t}-{\bf{R}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t}))$, and $\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}$, we have
@@
$\quad\leq 3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (56)
Combining (50) and (V), we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{2(1-p)}{\chi^{2}}\left(3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}\right).$
Therefore, combining it and (V), the inequality (31) follows.
@@
Then, we consider the bound of $-\frac{2\alpha}{n}\sum_{i=1}^{n}\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{e}}}^{t}\rangle$. Since $f_{i}$ is $L$-smooth and $\mu$-strongly convex, by (14), it gives that
@@
Substituting (V) and (V) into (V), and using $f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})\geq 0$, we have
@@ VI Proof of Theorem 3
$\alpha\leq\sqrt{\frac{(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L}\Longrightarrow\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\leq 3.$
@@
$\tilde{\gamma}$ $=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2(1-p)\big{(}3+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\big{)}}{\chi^{2}}$
$\leq\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}$
$\leq\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}.$
@@
$\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{12(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{2}.$
@@
$\frac{32\alpha^{2}L^{2}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{12(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{6},$
and using $\gamma=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}$, we have
$\alpha\leq\min\left\{\frac{1-\lambda_{2}}{32\sqrt{3}\chi L},\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}.$
@@
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{1-\lambda_{2}}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{(1-\lambda_{2})^{3}}\text{ and }\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{1024\alpha^{4}L^{4}\chi^{3}}{(1-\lambda_{2})^{3}}.$
@@
$\alpha\leq\sqrt[4]{\frac{(1-\lambda_{2})^{3}}{8\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}.$
@@
$+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\Big{(}\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}}\Big{)}$
$=$ $f(\bar{{\bf{x}}}^{t})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\frac{\alpha}{2}\Big{(}1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}\Big{)}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}$
$+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}^{t}-\frac{\alpha}{4}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}},$
where the last inequality holds because the condition (34) implies $\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{(1-\tilde{\gamma})(1-\gamma)}$. Taking full expectation, we have
$\mathbb{E}\!\left[\mathcal{L}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}^{t}\right]-\frac{\alpha}{4}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\frac{\alpha}{4}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq\mathcal{L}^{0}+T\Big{(}\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big{)},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]\leq\frac{4\mathcal{L}^{0}}{\alpha T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{16\sigma^{2}L^{4}\alpha^{4}\frac{2\chi}{1-\lambda_{2}}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{16L^{2}\sigma^{2}\alpha^{2}(2\chi^{2}+(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\leq f(\bar{{\bf{x}}}^{0})-f^{\star}+\frac{32\chi^{2}\alpha^{3}L^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}}.$ (61)
@@
$\frac{1}{1-\gamma}\leq\frac{4\chi}{1-\lambda_{2}},\ \frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}+\frac{128\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+(1-p))}{(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{1-\lambda_{2}}{32\sqrt{3}\chi L}$, we have $\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{n(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{3n(1-\lambda_{2})}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{2(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\chi^{3}+256\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+(1-p))}{2(1-\lambda_{2})\chi^{2}},$

### v4

@@ X Proof of Lemma 3
$\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\leq\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}}{1-\gamma}\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}-2\alpha\langle{\bf{\Gamma}}\mathcal{E}^{t},{\bf{C}}{\bf{S}}^{t}\rangle+4\alpha^{2}\|{\bf{S}}^{t}\|_{\mathrm{F}}^{2}.$
@@
${\bf{Q}}^{-1}={\bf{Q}}_{1}\left[\begin{array}[]{cc}\frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \frac{1}{2}\hat{{\bf{\Lambda}}}_{a}^{-\frac{1}{2}}&-\frac{j}{2}({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-\frac{1}{2}}\\ \end{array}\right]\text{ and }{\bf{I}}-\hat{{\bf{\Lambda}}}_{a}=\frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}.$
@@
By $\|({\bf{I}}-\hat{{\bf{\Lambda}}}_{a})^{-1}\|=\frac{2\chi}{p(1-\lambda_{2})}$ and $L$-smoothness of $f_{i}$, we have
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 4L^{2}\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{p(1-\lambda_{2})}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (53)
@@
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq 16L^{2}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{2\chi nL^{2}}{p(1-\lambda_{2})}\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{t}-\bar{{\bf{x}}}^{t+1}\|^{2}\;|\;\mathcal{G}^{t}\right].$ (54)
@@
$\mathbb{E}\!\left[\|{\bf{Q}}^{-1}{\bf{F}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $(16L^{2}+\frac{16\alpha^{2}L^{4}\chi}{p(1-\lambda_{2})})\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{2}L^{2}\chi}{p(1-\lambda_{2})}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{2}L^{2}\sigma^{2}\chi}{p(1-\lambda_{2})}.$ (55)
Thus, combining (X) and (55), it holds that
$\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=$ $\gamma\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$
$+\frac{8n\alpha^{4}L^{2}\chi}{p(1-\gamma)(1-\lambda_{2})}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{4\alpha^{4}L^{2}\sigma^{2}\chi}{p(1-\gamma)(1-\lambda_{2})}+4n\alpha^{2}\sigma^{2}.$ (56)
Then, we bound $\mathbb{E}\!\left[\|{\bf{Q}}^{-1}\hat{{\bf{P}}}^{\sf T}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$. Using $\|{\bf{X}}^{t}-\bar{{\bf{X}}}^{t}\|_{\mathrm{F}}^{2}\leq 4\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}$, $\hat{{\bf{Z}}}^{t}={\bf{X}}^{t}-{\bf{R}}^{t}-\alpha(\nabla F({\bf{X}}^{t})-\nabla F(\bar{{\bf{X}}}^{t})+{\bf{S}}^{t}))$, and $\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{p(1+\lambda_{n})(1-\lambda_{2})}$, we have
@@
$\quad\leq 3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{p(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}.$ (57)
Combining (51) and (X), we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{2p(1-p)}{\chi^{2}}\left(3\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{24\chi\alpha^{2}L^{2}}{p(1+\lambda_{n})(1-\lambda_{2})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+n\alpha^{2}\sigma^{2}\right).$
Therefore, combining it and (X), the inequality (32) follows.
@@
Then, we consider the bound of $\frac{2\alpha}{n}\sum_{i=1}^{n}\langle\nabla f_{i}({\bf{x}}_{i}^{t}),\bar{{\bf{e}}}^{t}\rangle$. Since $f_{i}$ is $L$-smooth and $\mu$-strongly convex, by (15), it gives that
@@
Substituting (X) and (X) into (X), and using $f(\bar{{\bf{x}}}^{t})-f({\bf{x}}^{\star})\geq 0$, we have
@@ XI Proof of Theorem 3
$\alpha\leq\sqrt{\frac{p(1+\lambda_{n})(1-\lambda_{2})}{2\chi}}\frac{1}{2L}\Longrightarrow\frac{24\chi\alpha^{2}L^{2}}{p(1+\lambda_{n})(1-\lambda_{2})}\leq 3.$
@@
$\tilde{\gamma}$ $=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2(1-p)p\big(3+\frac{24\chi\alpha^{2}L^{2}}{p(1+\lambda_{n})(1-\lambda_{2})}\big)}{\chi^{2}}$
$\leq\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{12p(1-p)}{\chi^{2}}$
@@
$\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{12p(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{2}.$
@@
$\frac{32\alpha^{2}L^{2}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\leq\frac{1-\gamma}{6},\quad\frac{12p(1-p)}{\chi^{2}}\leq\frac{1-\gamma}{6},$
and using $\gamma=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}$, we have
$\alpha\leq\min\left\{\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L},\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{12\chi^{3}}}\frac{1}{4L}\right\},\ \chi\geq\frac{288(1-p)}{1-\lambda_{2}}.$
@@
Since $\frac{1}{1-\gamma}\leq\frac{4\chi}{p(1-\lambda_{2})}$ and $\frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma}$, we have
$\frac{\frac{2\chi}{p(1-\lambda_{2})}}{(1-\gamma)^{2}}\leq\frac{32\chi^{3}}{p^{3}(1-\lambda_{2})^{3}}\text{ and }\frac{16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}\leq\frac{1024\alpha^{4}L^{4}\chi^{3}}{p^{3}(1-\lambda_{2})^{3}}.$
@@
$\alpha\leq\sqrt[4]{\frac{p^{3}(1-\lambda_{2})^{3}}{8\chi^{3}}}\frac{1}{4L}\Rightarrow\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}.$
@@
$+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\Big(\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{\chi^{2}}\Big)$
$=$ $f(\bar{{\bf{x}}}^{t})-f^{\star}+\frac{2\alpha L^{2}}{n(1-\tilde{\gamma})}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}-\frac{\alpha}{2}\Big(1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}\Big)\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}$
$+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}$
$\leq$ $\mathcal{L}^{t}-\frac{\alpha}{4}\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}},$
where the last inequality holds because the condition (35) implies $\frac{1}{2}<1-\frac{16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{(1-\tilde{\gamma})(1-\gamma)}$. Taking full expectation, we have
$\mathbb{E}\!\left[\mathcal{L}^{t+1}\right]\leq\mathbb{E}\!\left[\mathcal{L}^{t}\right]-\frac{\alpha}{4}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]+\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\frac{\alpha}{4}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]\leq\mathcal{L}^{0}+T\Big(\frac{L\alpha^{2}\sigma^{2}}{2n}+\frac{4\sigma^{2}L^{4}\alpha^{5}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{4L^{2}\sigma^{2}\alpha^{3}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}\Big),$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]\leq\frac{4\mathcal{L}^{0}}{\alpha T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{16\sigma^{2}L^{4}\alpha^{4}\frac{2\chi}{p(1-\lambda_{2})}}{n(1-\tilde{\gamma})(1-\gamma)}+\frac{16L^{2}\sigma^{2}\alpha^{2}(2\chi^{2}+p(1-p))}{(1-\tilde{\gamma})\chi^{2}}.$
@@
$\leq f(\bar{{\bf{x}}}^{0})-f^{\star}+\frac{32\chi^{2}\alpha^{3}L^{2}\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}}.$ (62)
@@
$\frac{1}{1-\gamma}\leq\frac{4\chi}{p(1-\lambda_{2})},\ \frac{1}{1-\tilde{\gamma}}\leq\frac{2}{1-\gamma},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{np^{3}(1-\lambda_{2})^{3}}+\frac{128\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{p(1-\lambda_{2})\chi^{2}}.$
Since $\alpha\leq\frac{p(1-\lambda_{2})}{32\sqrt{3}\chi L}$, we have $\frac{1024\sigma^{2}L^{4}\alpha^{4}\chi^{3}}{np^{3}(1-\lambda_{2})^{3}}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{3np(1-\lambda_{2})}\leq\frac{\alpha^{2}L^{2}\sigma^{2}\chi}{2p(1-\lambda_{2})}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\big\|\nabla f(\bar{{\bf{x}}}^{t})\big\|^{2}\right]$ $\leq\frac{4(f(\bar{{\bf{x}}}^{0})-f^{*})}{\alpha T}+\frac{128\chi^{2}L^{2}\alpha^{2}\varsigma^{2}_{0}}{p^{2}(1-\lambda_{2})^{2}T}+\frac{2L\alpha\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}\chi^{3}+256\chi\alpha^{2}L^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{2p(1-\lambda_{2})\chi^{2}},$